Shift report 20030501 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu May 01 2003 - 15:45:12 EDT

  • Next message: Apache: "Shift report 20030501 16:00-24:00"
    ------------------------------------------------------------
    Date: Thu May  1 15:45:12 2003, Shift: 08:00-16:00
    Supervisor: BB, with: 
    Log book#14 pages 88 - 89
    
    Shift Summary:
    ------------------------------------------------------------
    1. RHIC access. No beam.
    
    2. During access 2 small tubes (#7 and #8) were removed from the inner ring of the left BeamBeam array. Their radiators and glue joints will be replaced to try to increase their overall gain.
    
    3. Dana mounted a prototype slat (prototype for TOFW3) behind TOFW2.
    
    4. Big tube #32 in the right BB array is holding voltage at 2800 V.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu May 01 2003 - 15:45:46 EDT